OpenStudy (anonymous):

simplify ((x^(2n)+2x^n)/(x^(n+1)+2x) *((x^2-3x)/(x^(n+1)-3x^n))

OpenStudy (kinggeorge):

Try multiplying the numerators together, and then multiply the denominators together. If you compare them, I'll bet they look the same.

OpenStudy (anonymous):

i guess my real question is how do you mulitply something like x^2n * x^(n+1) ?

OpenStudy (kinggeorge):

As a general rule,\[x^a\cdot x^b=x^{a+b}\]So using your example,\[x^{2n}\cdot x^{n+1}=x^{2n+n+1}=x^{3n+1}\]

OpenStudy (anonymous):

thank you!!

OpenStudy (kinggeorge):

you're welcome